I need help in solving this equation - I came up with x=7 and got it wrong.
3[x-2] +2=17
Thanks

Found 2 solutions by dabanfield, richwmiller:
Answer by dabanfield(803)

3[x-2] +2=17
3*x - 3*2 + 2 = 17
3*x - 6 + 2 = 17
3*x = 17 + 6 - 2
3*x = 9
x = 3

Answer by richwmiller(17219)

x does equal 7
3[x-2] +2=17
3x-*3*2=15
3x-6=15
3x=21
x=7
so unless you wrote the problem incorrectly x does equal 7
The other tutor made errors in his calculations
3[x-2] +2=17
3*x - 3*2 + 2 = 17
3*x - 6 + 2 = 17
3*x = 17 + 6 - 2
17+6-2=23-2=21 not 9
so
3x=21
x=7